Q. (3/4) + (3/4 + 1/100) + (3/4 +2/100) + ........... (3/4 + 99/100)


Find the greatest integer function.. plz sumbody xplain every step briefly

Grade:11

2 Answers

Rahul Saxena
12 years ago

3/4=0.75

now, till 3/4+24/100,

the sum of the series will be 0,      as[0]=0

3/4+25/100=1      [1]=1,

3/4+99/100=1.73          [1.73]=1

then from 3/4+25/100_ _ _ _ _ _3/4+99/100=75

therefore the sum of series=75         ans

xyz xz
12 years ago

the greatest integer function is 4

because {4.425} = 4

the sum of 3/4 + 3/4+1/100................................ (3/4+99/100) = 4.425

 therfore {4.425} = 4